OpenStudy (anonymous):

Please help will give medal and fan Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18

OpenStudy (anonymous):

12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = -15 3y = 9

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@KingGeorge

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

@master50777 hey do u think u could help me with this one too

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

wait what